OpenStudy (anonymous):

what is (17)/(2)m 8m the are is ? ?

OpenStudy (dumbcow):

area = 17/2 * 8/1 = 17*4 = 68

OpenStudy (anonymous):

don't listen to whatever dumbcow posts. it's gonna be incorrect